15.
Find the slope and y-intercept.
10x – 7y = 7

Answer by venugopalramana(3286)

Find the slope and y-intercept.
10x – 7y = 7
PUT IT IN STANDARD FORM OF
Y=MX+C..WHERE..M=SLOPE AND C=Y INTERCEPT
SO WE GET
7Y=10X-7
Y=10X/7 - 1
SLOPE =10/7
Y INTERCEPT = - 1